EPS-HEP99 Abstract # 000 Parallel sessions: Plenary sessions:

Determination of the Strong Coupling Constant from Inclusive Jet Cross Sections in Deep-Inelastic Positron Proton Collisions at HERA

H1 Collaboration,

DESY, Notkestr.85, D-22607 Hamburg, Germany

We have measured the inclusive jet cross section in the Breit frame in deep-inelastic scattering at large momentum transfers Q^{2} and large transverse jet energies E_{T}. The analysis is based on data corresponding to integrated luminosity of L = 33 pb^{−1} taken in the years 1995-1997 with the H1 detector at the ep collider HERA at \sqrt{s}_{ep} = 300 GeV. The inclusive jet cross section is presented double differentially as a function of both hard scales for 50 < E_{T}^{2} < 2500 GeV^{2} and 150 < Q^{2} < 5000 GeV^{2}. In a QCD analysis of the data we have determined the strong coupling constant α_{s}. To investigate the energy dependence of α_{s} we extract values of α_{s}(μ_{r}) from subsets of the data in four regions of the renormalization scale which is either chosen to be μ_{r} = E_{T} or μ_{r} = Q. Finally we have performed a combined fit to the data over the whole range of E_{T}^{2} and Q^{2} and obtain a combined result of α_{s}(M_{Z}).